OpenStudy (anonymous):

Ohm’s Law is represented by the equation I=V/R. Explain how the current would change if the amount of resistance decreased and the voltage stayed the same. (HINT: try plugging numbers into the equation for R, but make sure V stays the same. Then describe what happens to I(current)

OpenStudy (anonymous):

someone please help!! medal, fan!

OpenStudy (anas.p):

K listen carefully, this might get a little tricky. We will follow the hint for now. I=V/R Lets say V=20 Volts and R=4 ohms we will get I=20/4 which implies I=5 A. Now we will keep Voltage constant, and decrease the value of R from 4ohms to 2ohms. we will now get I=20/2. which implies I=10A. So as we decreased the value of the R, we got an increased value of I, provided V is kept constant. Hope that helped. Any doubts? please ask.

OpenStudy (anonymous):

thanks for that one!! i understand it. could you help me with anouther @anas.p